# RoundTable: Investigating Group Decision-Making Mechanism in Multi-Agent Collaboration

### v1

@@ A.2 RoundTable Prompt Design
# Collaboration Rules You start with nothing decided. The intermediate result will be decided by the social choice function at the end of each round. In each round, the collaboration runs in 3 phases with the following order: 1. Message Phase: At the beginning of each round, you can send one message to a shared channel for either Talking to one or more agents. All agents will send messages simultaneously. You will be able to see all messages from all agents after the end of the message phase. 2. Proposal Phase: After the end of the message phase, you will have the opportunity to propose potential solution. If you don’t propose in this phase, your latest proposal will be used for voting. 3. Voting Phase: At the end of the round, all agents’ latest proposal will be voted. When agents didn’t propose in this round, their latest proposal will be used for voting. All votes will be processed with the social choice function: name_of_social_choice, where explanation_of_social_choice. If the social choice function selects a proposal, the intermediate result will be updated accordingly. After each round, each agent will be able to see the result of the vote from the previous round and the conversation history from all rounds. The collaboration will run for max_rounds rounds. After the last round, the latest result will be the final result. # Your Background my_agent_backround
@@
Message Phase Prompt You are my_name, currently in the message phase of round round_num. In this phase, you can: 1. Answer questions posed by others. 2. Share your findings or insights. 3. Ask questions to further the discussion. You may engage in multiple activities using multiple sentences. Please type your message in the following JSON format: {“target”: <list of agent names>, “message”: <str, your message>} Don’t generate anything except the JSON format.
@@
Proposal Phase Prompt You are my_name, currently in the proposal phase of round round_num. You have an opportunity to make a proposal of the potential solution. Whether or not you submit a new proposal, your latest proposal will be considered as a candidate proposal for the voting phase. You have two options: 1. Make a proposal: - You can propose a potential solution by the provided format. 2. Do not make a proposal: - If you do not want to propose a solution, you can return None as your proposal. Please type your proposal in the following JSON format: {“reason_for_decision”: <your step by step reasoning for your decision>, “proposal”: proposal_format_text, or None} Don’t generate anything except the JSON format.
@@
Voting Phase Prompt (for Vote-Based Mechanisms) You are my_name, at the voting phase at round round_num. In this phase, all agents’ latest proposal will be voted by name_of_social_choice, where explanation_of_social_choice. If the social choice function selects a proposal, the intermediate result will be updated accordingly. You have two actions to choose: vote or not vote. 1. For vote: - You can only vote for one of the proposals from the candidate list. 2. For not vote: - You should vote None. - If you do not want to vote for any of the proposals, you can vote None. - If there is no proposal, you vote None. The same proposal proposed by multiple agents will be merged as one proposal. If there are multiple proposals passed, none of the proposals will be selected. If no proposals are passed, the current intermediate result will be kept. The current candidate proposals are as follows: proposal_list
@@
What is your vote? Please answer in the following JSON format: {“reason_for_decision”: <your step by step reasoning for your decision>, “decision”: <id of the proposal from the candidates you want to vote, or None>} Don’t generate anything except the JSON format.
Appendix B Details of Social Choice Methods
We present the definition and the prompt we use to define different social choice methods.
@@ Appendix B Details of Social Choice Methods
Ranked Voting Description Each agent ranks all candidate proposals from the most preferred to the least preferred. Social Choice will assign 1, 1/2, 1/3… points to the 1st, 2nd, 3rd… candidates on each ballot. The proposal with the highest total points will be selected.
@@
Appendix C Details of Experiments
C.1 Simple and Complex Environments
Simulated environments are simplified, abstract representations of the world. They include economic experiments, such as those based on game theory, and games like chess. The primary advantages of simulated environments are: 1. They allow for the emphasis on specific agent behaviors within a controlled setting by modifying the environmental design; 2. The utility function can be mathematically expressed, and the quality of each proposal can be quantified numerically.
Complex environments are real-world applications that are more intricate and challenging, offering a closer reflection of scenarios where agents must handle numerous variables and unpredictability. These include tasks like collaborative problem-solving, negotiations, and real-time strategy games. While evaluation in these settings can be subjective, they provide deeper insights into how agents adapt, cooperate, and make decisions in dynamic, less controlled conditions, highlighting their resilience and versatility.
C.2 Exchange Economy
@@
The task description we use is as follows:
@@
Below is the format of an agent’s goal:
Exchange Economy Agent Goal Your goal is to maximize your individual utility function by communicating, proposing, and voting with other agents. Your utility function is util_func
C.3 Exchange Economy: Other Utility Sets
@@
C.4 Recommendation System
Recommendation system aims to reveal unique collaborative behaviors that exchange economic does not have. First, there exists strong information asymmetry between agents. Any piece of information is not sufficient to correctly predict the rating. Secondly, the group decision-making in these systems encourages diverse approaches to problem-solving. Agents are not only consumers of information but also contributors, creating a feedback loop where the quality of recommendations improves with increased participation and collaboration. Lastly, the value generated in such a system is not solely based on the final recommendation but also on the process of reaching that recommendation. The interactions, negotiations, and information exchanges that occur along the way contribute to the overall effectiveness and satisfaction of the system.
Below is the description of the environment sent to agents:
@@ C.4 Recommendation System
In the recommendation system enviroment, all agents share the same goal, but has different background dataset. Here is the goal used in our setting:
Recommendation System Agent Goal Your goal is to predict the rating the target user would give to the target movie. Utilize all available information about both the user and the movie to make the most accurate prediction possible. You only have access to data_access, but you can communicate with other agents to get more information. # Your Data: agent_dataset
@@
1 Toy Story [’Animation’, ”Children’s”, ’Comedy’] 19950101 3 19980331
@@ 50 Star Wars [’Action’, ’Adventure’, ’Romance’, ’Sci-Fi’, ’War’] 19770101 4 19980331
C.5 Dialogue Act Labeling
@@
C.6 Dialogue Act Transition Graph
@@
$p_{A\rightarrow B}=\frac{|A\rightarrow B|}{|A|}=\frac{\sum_{s}\sum_{r=1}^{10}|\{(i,\neg i)|i\in{\mathbb{I}},A\in{\mathbb{D}}{\mathbb{A}}_{i,r-1},B\in{\mathbb{D}}{\mathbb{A}}_{\neg i,r}\}|}{\sum_{s}\sum_{r=1}^{10}|\{i|i\in{\mathbb{I}},A\in{\mathbb{D}}{\mathbb{A}}_{i,r-1}\}|}$ (1)
Where $s$ is the index of simulations, ${\mathbb{D}}{\mathbb{A}}_{i,r}$ is a set of dialogue acts from the message of the agent $i$ at round $r$, and ${\mathbb{D}}{\mathbb{A}}_{r=0}=\{Start\},{\mathbb{D}}{\mathbb{A}}_{r=11}=\{End\}$.
C.7 Early Stopping in MAS
@@
We examine all combinations of hyperparameters on the training data, and identifies the best sets for each social choice and environment. Using the best hyperparameters, we evaluate the performance of early stopping on the test set.
C.8 Models
For all experiments, we by default use gpt-4o-mini-2024-07-18 (OpenAI, 2023) as the main LLM for agents. In ablation study, we use gpt-3.5-turbo-0125, gpt-4o-2024-05-13, Llama-3.1-70b-Instruct and Llama-3.1-8b-Instruct for comparison (Dubey et al., 2024). To calculate sentence embeddings we used paraphrase-MiniLM-L6-v2 (Reimers & Gurevych, 2019). For dialogue act labeling, we use Llama-3.1-8b-Instruct. For all LLM inferences, we used temperature = 0.
Appendix D Details of Results
D.1 Ablation Studies
@@
AUC@3 AUC@5 AUC@10 RATIONALITY MIN/MAX RIGIDITY
Social Choice: Majority, # Agent: $\mathbf{K=3}$
@@
Social Choice: Majority, Model: gpt-4o-mini
@@
The result of ablation studies is shown in Table 6. We further compare the performance of multi-agent collaboration with different LLMs in Figure 2(b). For simplicity, we only compare $K=3$ agent system in majority voting setting with 30 simulations. The results of gpt-3.5-turbo-0125, gpt-4o-2024-05-13, Llama-3.1-8b-Instruct and Llama-3.1-70b-Instruct in Table 1 show that multi-agent collaboration works with different LLMs. The stronger (larger) the model, the better performance observed in collaboration, both in end round performance and also efficiency in reaching to an agreement. Notably, the Rationality of the weaker model, gpt-3.5-turbo-0125 and Llama-3.1-8b-Instruct, is only 17.88% and 3.78%, indicates that single agent will largely fail due to the complexity of the task. Nonetheless, the final round performance has reached 64% and 80%, shows the effectiveness of multi-agent collaboration with weak models.
Scaling on MAS is evaluated with different number of agents. From the nature of decentralized collaboration, the more the agents, the harder to reach an agreement. Furthermore, the additional amount of historical context brought by increasing number of agents can also be a burden to the agent’s performance. Comparing $K=3,4,5$ settings in Table 1, the quality and efficiency of collaboration generally drops with the increasing number of agents. Exceptionally, performance in $K=4$ setting is worse than that of $K=5$. This is because majority voting is harder to achieve an agreement with, even number of participants. With 4 agents, a proposal needs at least 3 votes to get accepted; 5 agent setting also requires 3 votes, but they can ignore up to 2 agents’ preferences or agreements.
D.2 Recommendation System
We show the detailed result between social choices in recommendation system environment in Table 7.
Table 7: Recommendation system environment results comparison between social choices. The smallest and largest value in a category is colored in blue and red. Experiments are made with gpt-4o-mini.
@@ D.2 Recommendation System
D.3 Cross-Agent Conversation Analysis
D.3.1 Linguistic Statistics
In this section, we show how linguistic features presented in different environments and social choices. Table 8 is heatmaps for average message length, message complexity, and information difference observed in agent conversation.
Table 8: Heatmaps of average message length, message complexity, and information difference in each round. Analysis is made in 3 agents, gpt-4o-mini setting. Color gradient is calculated with green as maximum, red as minimum, and white as median value in each table.

### v2

@@ B.2 RoundTable Prompt Design
# Collaboration Rules You start with nothing decided. The intermediate result will be decided by the voting rule at the end of each round. In each round, the collaboration runs in 3 phases with the following order: 1. Message Phase: At the beginning of each round, you can send one message to a shared channel for either Talking to one or more agents. All agents will send messages simultaneously. You will be able to see all messages from all agents after the end of the message phase. 2. Proposal Phase: After the end of the message phase, you will have the opportunity to propose potential solution. If you don’t propose in this phase, your latest proposal will be used for voting. 3. Voting Phase: At the end of the round, all agents’ latest proposal will be voted. When agents didn’t propose in this round, their latest proposal will be used for voting. All votes will be processed with the voting rule: name_of_voting_rule, where explanation_of_voting_rule. If the voting rule selects a proposal, the intermediate result will be updated accordingly. After each round, each agent will be able to see the result of the vote from the previous round and the conversation history from all rounds. The collaboration will run for max_rounds rounds. After the last round, the latest result will be the final result. # Your Background my_agent_backround
@@
Figure 6: Prompt used for initialization stage in RoundTable. Italicized words represent inputs or variables within the prompt.
@@
Message Phase Prompt You are my_name, currently in the message phase of round round_num. In this phase, you can: 1. Answer questions posed by others. 2. Share your findings or insights. 3. Ask questions to further the discussion. You may engage in multiple activities using multiple sentences. Please type your message in the following JSON format: {“target": <list of agent names>, “message": <str, your message>} Don’t generate anything except the JSON format.
Figure 7: Prompt used for message phase in RoundTable. Italicized words represent inputs or variables within the prompt.
@@
Proposal Phase Prompt You are my_name, currently in the proposal phase of round round_num. You have an opportunity to make a proposal of the potential solution. Whether or not you submit a new proposal, your latest proposal will be considered as a candidate proposal for the voting phase. You have two options: 1. Make a proposal: - You can propose a potential solution by the provided format. 2. Do not make a proposal: - If you do not want to propose a solution, you can return None as your proposal. Please type your proposal in the following JSON format: {“reason_for_decision": <your step by step reasoning for your decision>, “proposal": proposal_format_text, or None} Don’t generate anything except the JSON format.
Figure 8: Prompt used for proposal phase in RoundTable. Italicized words represent inputs or variables within the prompt.
@@
Voting Phase Prompt (for Discrete Voting Rules) You are my_name, at the voting phase at round round_num. In this phase, all agents’ latest proposal will be voted by name_of_social_choice, where explanation_of_social_choice. If the voting rule selects a proposal, the intermediate result will be updated accordingly. You have two actions to choose: vote or not vote. 1. For vote: - You can only vote for one of the proposals from the candidate list. 2. For not vote: - You should vote None. - If you do not want to vote for any of the proposals, you can vote None. - If there is no proposal, you vote None. The same proposal proposed by multiple agents will be merged as one proposal. If there are multiple proposals passed, none of the proposals will be selected. If no proposals are passed, the current intermediate result will be kept. The current candidate proposals are as follows: proposal_list
@@
What is your vote? Please answer in the following JSON format: {“reason_for_decision": <your step by step reasoning for your decision>, “decision": <id of the proposal from the candidates you want to vote, or None>} Don’t generate anything except the JSON format.
Figure 9: Prompt used for voting phase in RoundTable. Italicized words represent inputs or variables within the prompt.
Appendix C Details of Voting Rules
We present the definition and the prompt we use to define different voting rules, including unanimous voting (Figure 10), majority voting (Figure 11), plurality voting (Figure 12), rated voting (Figure 13), ranked voting (Figure 14), and cumulative voting (Figure 15). The formal definition of voting rules are as below:
Unanimous Voting (Arrow, 2012): The proposal $x^{*}_{r}$ that receives votes from all agents is selected:
$\hat{x^{*}}_{r}=\arg\max_{c\in C}\mathbbm{1}\left(\sum_{i\in\mathcal{I}}v_{i,c,r}=|\mathcal{I}|\right)$
Majority Voting (Arrow, 2012): The candidate $c$ that receives votes from more than half of all agents is selected:
$\hat{x^{*}}_{r}=\arg\max_{c\in C}\mathbbm{1}\left(\sum_{i\in\mathcal{I}}v_{i,c,r}>\frac{|\mathcal{I}|}{2}\right)$
Plurality Voting (Arrow, 2012): The candidate $c$ that receives the most votes is selected:
$\hat{x^{*}}_{r}=\arg\max_{c\in C}\sum_{i\in\mathcal{I}}v_{i,c,r}$
Rated Voting (Baujard et al., 2018): Each agent rates proposals on a 5-point scale, and the candidate $c$ with the highest total score is selected:
$\hat{x^{*}}_{r}=\arg\max_{c\in C}\sum_{i\in\mathcal{I}}s_{i,c,r},\quad s_{i,c,r}\in\{1,2,3,4,5\}$
Ranked Voting (Arrow, 2012): Agents rank proposals by preference, assigning points as $1,1/2,1/3,$ etc., and the candidate $c$ with the highest total score is selected:
$\hat{x^{*}}_{r}=\arg\max_{c\in C}\sum_{i\in\mathcal{I}}\frac{1}{rank_{i,c,r}}$
where $rank_{i,c,r}$ is the rank of proposal $c$ given by agent $i$ (lower rank values indicate higher preference).
Cumulative Voting (Black, 1958): Each agent distributes $|\mathcal{I}|$ points among $|\mathcal{I}|$ proposals, and the candidate $c$ with the highest total is selected:
$\hat{x^{*}}_{r}=\arg\max_{c\in C}\sum_{i\in\mathcal{I}}pt_{i,c,r}$
where $pt_{i,c,r}$ represents the number of points agent $i$ assigns to proposal $c$ such that:
$\sum_{c\in C}p_{i,c,r}=|\mathcal{I}|$
@@ Appendix C Details of Voting Rules
Figure 10: Prompt used for describing unanimous voting in RoundTable.
@@
Figure 11: Prompt used for describing majority voting in RoundTable.
@@
Figure 12: Prompt used for describing plurality voting in RoundTable.
@@
Figure 13: Prompt used for describing rated voting in RoundTable.
Ranked Voting Description Each agent ranks all candidate proposals from the most preferred to the least preferred. 1, 1/2, 1/3… points will be assigned to the 1st, 2nd, 3rd… candidates on each ballot. The proposal with the highest total points will be selected.
Figure 14: Prompt used for describing ranked voting in RoundTable.
@@
Figure 15: Prompt used for describing cumulative voting in RoundTable.
Appendix D Details of Experiments
D.1 Exchange Economy
@@
The task description we use is presented in Figure 16. Prompt for agent’s objective is shown in Figure 17.
@@
Figure 16: Prompt used for describing the task in exchange economy environment in RoundTable. Italicized words represent inputs or variables within the prompt.
Exchange Economy Agent Objective Your goal is to maximize your individual utility function by communicating, proposing, and voting with other agents. Your utility function is util_func
Figure 17: Prompt used for describing an agent’s objective in exchange economy environment in RoundTable. Italicized words represent inputs or variables within the prompt.
D.2 Exchange Economy: Other Utility Sets
@@
D.3 Recommendation System
Recommendation system aims to reveal unique collaborative behaviors that exchange economic does not have. First, there exists strong information asymmetry between agents. Any piece of information is not sufficient to correctly predict the rating. Secondly, the group decision-making in these systems encourages diverse approaches to problem-solving. Agents are not only consumers of information but also contributors, creating a feedback loop where the quality of recommendations improves with increased participation and collaboration. Lastly, the value generated in such a system is not solely based on the final recommendation but also on the process of reaching that recommendation. The interactions, negotiations, and information exchanges that occur along the way contribute to the overall effectiveness and satisfaction of the system. We show the description of the environment in Figure 18. In the recommendation system enviroment, all agents share the same objective, but has different background dataset. The prompt used for an agent’s objective is in Figure 19.
@@ D.3 Recommendation System
Figure 18: Prompt used for describing the task in recommendation environment in RoundTable. Italicized words represent inputs or variables within the prompt.
Recommendation System Agent Objective Your goal is to predict the rating the target user would give to the target movie. Utilize all available information about both the user and the movie to make the most accurate prediction possible. You only have access to data_access, but you can communicate with other agents to get more information. # Your Data: agent_dataset
Figure 19: Prompt used for describing an agent’s objective in recommendation environment in RoundTable. Italicized words represent inputs or variables within the prompt.
@@
1 Toy Story [’Animation’, "Children’s", ’Comedy’] 19950101 3 19980331
@@ 50 Star Wars [’Action’, ’Adventure’, ’Romance’, ’Sci-Fi’, ’War’] 19770101 4 19980331
D.4 Language Features
Information Distance measures how much new information is introduced. Effective collaboration should consistently bring new insights, while low information gain signals a stalled conversation. It is calculated as the average cosine distance between messages in the current round and the center embedding of the previous round.
Dialogue Acts are communicative functions that capture actions, intents, and behaviors within messages as discrete states. We design a set of dialogue acts for multi-agent collaboration: Inform, Request, Confirm, Summarize, Evaluate, Propose, Compromise, Defend, Accept, Decline, and use LLM-labeling to perform multi-label classification on a target message based on the previous round. Details of the dialogue act labels and prompt can be found in Appendix D.5.
D.5 Dialogue Act Labeling
@@
D.6 Dialogue Act Transition Graph
We constructed transition graphs to visualize the dynamics of dialogue act transitions across rounds, highlighting the most frequent transitions. Each node represents a dialogue act, and directed edges indicate the probability of transitions between acts from one round to the next. The figure displays only the most probable edge per node, excluding self-loops.
@@ D.6 Dialogue Act Transition Graph
$p_{A\rightarrow B}=\frac{|A\rightarrow B|}{|A|}=\frac{\sum_{s}\sum_{r=1}^{10}|\{(i,\neg i)|i\in\mathcal{I},A\in\mathcal{D}_{i,r-1},B\in\mathcal{D}_{\neg i,r}\}|}{\sum_{s}\sum_{r=1}^{10}|\{i|i\in\mathcal{I},A\in\mathcal{D}_{i,r-1}\}|}$
Where $s$ is the index of a collaboration, $\mathcal{D}_{i,r}$ is a set of dialogue acts from the message of the agent $i$ at round $r$, and $\mathcal{D}_{r=0}=\{Start\},\mathcal{D}_{r=11}=\{End\}$.
D.7 Early Stopping in MAS
This section explains the detailed definitions of early stopping methods. In this experiment, we evaluate the following early stopping methods:
First Agreement stops collaboration whenever a proposal passes the voting threshold.
Information Distance is the average embedding distance for messages between adjacent rounds captured at the round that produced the best outcomes in the train set, and stops collaborations in the test set when the distance became lower than the threshold. This idea is supported by the observation in Section 5.
Dialogue Act method utilizes pairs of dialogue acts, linking one from the previous round to another from the current round. We perform ordinary least squares (OLS) regression on all such pairs in relation to the performance. The regression coefficients indicate the most impactful pairs that act as stop signals. Further details of the algorithm are provided in Appendix D.7.1.
D.7.1 Dialogue Act Early Stopping Method
@@
We examine all combinations of hyperparameters on the training data, and identifies the best sets for each voting rule and environment. Using the best hyperparameters, we evaluate the performance of early stopping on the test set.
D.8 Models and Dataset
For all experiments, we by default use gpt-4o-mini-2024-07-18 (OpenAI, 2023) as the main LLM for agents. In ablation study, we use gpt-3.5-turbo-0125, gpt-4o-2024-05-13, Llama-3.1-70b-Instruct and Llama-3.1-8b-Instruct for comparison (Dubey et al., 2024). To calculate sentence embeddings we used paraphrase-MiniLM-L6-v2 (Reimers and Gurevych, 2019). For dialogue act labeling, we use Llama-3.1-8b-Instruct. For all LLM inferences, we used temperature = 0.
MovieLens 100K dataset is used in recommendation system, which is under CC0: Public Domain license Harper and Konstan (2015). Our use of the dataset is consistent with the intended use. MovieLens does not contain personally identifying info or offensive content.
Appendix E Details of Results
Figure 20: LLM comparison with group total utility in exchange economy, using 3 agent, asymmetric, majority voting setting. Line plots (left y-axis) show the group total utility achieved over rounds; bar plots (right y-axis) represent the ratio of cases where participants yet to reach an agreement until a certain round. The red horizontal line indicates the maximum achievable group total utility.
Figure 21: Comparison between voting rules in recommendation system environment. Line plots (left y-axis) show the comparison of RMSE achieved over rounds; bar plots (right y-axis) represent the ratio of cases where participants yet to reach an agreement until a certain round. The green horizontal line is baseline (Always Guess 4), and the blue line is SoTA performance Behera and Nain (2023).
E.1 Ablation Studies
@@
AUC@3 AUC@5 AUC@10 RATIONALITY FAIRNESS RIGIDITY
Voting Rule: Majority, # Agent: $\mathbf{K=3}$
@@
Voting Rule: Majority, Model: gpt-4o-mini
@@
The result of ablation studies is shown in Table 6. We further compare the performance of multi-agent collaboration with different LLMs in Figure 20. For simplicity, we only compare $K=3$ agent system in majority voting setting with 30 collaborations. The results of gpt-3.5-turbo-0125, gpt-4o-2024-05-13, Llama-3.1-8b-Instruct and Llama-3.1-70b-Instruct in Table 1 show that multi-agent collaboration works with different LLMs. The stronger (larger) the model, the better performance observed in collaboration, both in end round performance and also efficiency in reaching to an agreement. Notably, the Rationality of the weaker model, gpt-3.5-turbo-0125 and Llama-3.1-8b-Instruct, is only 17.88% and 3.78%, indicates that single agent will largely fail due to the complexity of the task. Nonetheless, the final round performance has reached 64% and 80%, shows the effectiveness of multi-agent collaboration with weak models.
Scaling on MAS is evaluated with different number of agents. Due to the nature of decentralized collaboration, the more the agents, the harder to reach an agreement. Furthermore, the additional amount of historical context brought by increasing number of agents can also be a burden to the agent’s performance. Comparing $K=3,4,5$ settings in Table 1, the quality and efficiency of collaboration generally drops with the increasing number of agents. Exceptionally, performance in $K=4$ setting is worse than that of $K=5$. This is because majority voting is harder to achieve an agreement with, even number of participants. With 4 agents, a proposal needs at least 3 votes to get accepted; 5 agent setting also requires 3 votes, but they can ignore up to 2 agents’ preferences or agreements.
E.2 Recommendation System
We show the detailed result between voting rules in the recommendation system environment in Table 7.
@@ E.2 Recommendation System
Table 7: Recommendation system environment results comparison between voting rules. The smallest and largest value in a category is colored in blue and red. Experiments are made with gpt-4o-mini.
E.3 Cross-Agent Communication Analysis
In this section, we show how language features are presented in different environments and voting rules. Table 8 is heatmaps for average message length, message complexity, and information distance observed in agent conversation. Message length increases over time in both environments, with shorter messages in early rounds and longer ones later. The recommendation system generally has longer messages due to its need for detailed explanations. Message complexity in the exchange economy starts low, peaks early, dips, and then gradually increases toward the final round, while the recommendation system shows consistently higher and increasing complexity throughout. Information distance steadily decreases in both environments, indicating a convergence of topics and less novel information as the discussions progress.